OpenStudy (anonymous):

Help? x - 0.1x + 0.07(x - 0.1x) = 69.62

OpenStudy (anonymous):

distribute the 0.07 to the x and -0.1x

OpenStudy (anonymous):

x - 0.1x + 0.07x - .007x = 69.62

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

pretty sure I can

OpenStudy (anonymous):

0.9x+.07(0.9x)=69.62 (1.7)(.9)x=69.62 x=69.62/(.9)(1.7) x=69.62/1.53 x=45.5032